Carly is five less than twice John's age. If the sum of their ages is 25 find both of their ages
Answer by ankor@dixie-net.com(22740)

let c = C's age
let j = J's age
:
write an equation for each statement
:
Carly is five less than twice John's age.
c = 2j - 5
If the sum of their ages is 25
c + j = 25
find both of their ages
Replace c in the above equation with (2j-5), from the first statement
(2j-5) + j = 25
2j + j = 25 + 5
3j = 30
j = 30/3
j = 10 yrs is J's age
I'll let you find C's age by substituting 10 for j in the first equation
Check your answers in the 2nd statement